Hello, Can you please help with this problem and provide explanation for the correct answer? Thank...

Hello,

Can you please help with this problem and provide explanation for the correct answer? Thank you

Harold Reese must choose between two bonds:

Bond X pays $70 annual interest and has a market value of $845. It has 10 years to maturity.
Bond Z pays $60 annual interest and has a market value of $870. It has five years to maturity.

Assume the par value of the bonds is $1,000.

a. Compute the current yield on both bonds. (Do not round intermediate calculations. Input your answers as a percent rounded to 2 decimal places.)

b. Which bond should he select based on your answers to part a?

  • Bond X

  • Bond Z

c. A drawback of current yield is that it does not consider the total life of the bond. For example, the approximate yield to maturity on Bond X is 9.43 percent. What is the approximate yield to maturity on Bond Z? The exact yield to maturity? (Use the approximation formula to compute the approximate yield to maturity and use the calculator method to compute the exact yield to maturity. Do not round intermediate calculations. Input your answers as a percent rounded to 2 decimal places.)

d. Has your answer changed between parts b and c of this question?

  • Yes

  • No
Answer #1

Current yield gives information about rate of return that bond generates in a year. So formula is

Current yield = Annual interest amount / market price of bond

a) Current yield on bond X = $70 / $845 *100 = .0828*100= 8.28%

Current yield on bond Z = $60 / $870*100 = .0689*100= 6.89%

b) Harold Reese should select bond X as current yield of bond X is 8.28% which is higher than yield of bond Z i.e. 6.89%.

C) Approx yield to maturity (YTM) = C+ F-P/n / (F+P)/2 where

C = Interest amount, F = Face Value , P= Price, n = no of years to maturity

So approx YTM on bond Z = 60 + 1000-870/5 / (1000+870)/2

60+ (130/5) / 1870/2

60+26/935

9.19%

The exact yield to maturity on bond Z is 9.37%

d) the answer to d is No as still Bond X gives more returns than Bond Z.